OpenStudy (anonymous):

f(x)=x(sqrt(x^2+3))+10 Find the derivative, f'(1) and second deriv. f"(1) Then find the value of x where the second derivative equals 0. That is, solve the equation f"(x)=0 for x.

OpenStudy (campbell_st):

lots of ways ot do this rewriting is the 1st step \[f(x) = (x^4 + 3x^2)^{1/2} + 10\] use the chain rule the constant 10 will disappear put the x inside the square root \[u = x^4 + 3x^2\] \[du/dx = 4x^3 + 6x\] then \[y = u^{1/2}\] \[dy/du = 1/2u^{-1/2}\] \[dy/dx = dy/du \times du/dx\] then \[dy/dx = x \sqrt{x^2 + 3} \times (4x^3 + 6x)\] so\[f'(x) = (4x^4 + 6x^2)\sqrt{x^2 + 3}\] for f"(x) use the product rule... good luck